OpenStudy (anonymous):

Auditions for a play in New York were held with 120 men and women trying out for the lead singer position. If twice the number of men less three times the number of women is equal to 20, how many men auditioned? 44 40 76 80

OpenStudy (anonymous):

twice the number of men less three times the number of women is equal to 20 => 2(number of men) - 3(number of women) = 20 ....(1) also number of men + number of women = 120 ....(2) from the above two equations 2(number of men) - 3(120 - number of men) = 20 => 5(number of men) = 20 + 360 => number of men = 380/5 = 76